OpenStudy (anonymous):

How do I find the arguement of z, when z=(1+(sqrt3)i)^m?

OpenStudy (anonymous):

So like this: \[z=(1+(\sqrt3)i)^m\]

OpenStudy (anonymous):

@Loser66 ?

OpenStudy (loser66):

arg (z) = arctan \(\dfrac{y}{x} =arctan\dfrac{\sqrt3}{1}= arctan\sqrt3 = 60 ~~or ~~=\dfrac{\pi}{3}\)

OpenStudy (anonymous):

but what if the power is not an integer, but a variable m? I'm asked to provide the general formula, not just just for a specific case

OpenStudy (loser66):

oh yea, you are right.

OpenStudy (anonymous):

Do you have any other ideas?

OpenStudy (anonymous):

Because I'm pretty lost

OpenStudy (loser66):

I don't know, I am sorry. let tag other

OpenStudy (anonymous):

That's alright, don't worry about it :)

OpenStudy (asnaseer):

@Teinis -- do you know that any complex number can be written as:\[z=|z|e^{Arg(z)}\]

OpenStudy (anonymous):

Yes I'm familiar with that equation

OpenStudy (asnaseer):

ok, so we can then say that:\[z^m=|z|^m(e^{Arg(z)})^m=|z|^me^{mArg(z)}\]Therefore:\[Arg(z^m)=mArg(z)\]

OpenStudy (anonymous):

Although isn't it \[z=\left| z \right| e ^{iArgz}\]

OpenStudy (asnaseer):

yes - sorry I missed the i

OpenStudy (asnaseer):

Did you follow the argument above for \(z^m\)?

OpenStudy (anonymous):

I don't really get how you got from your first equation to your second. Can you explain it?

OpenStudy (asnaseer):